Journalist Report – May 16th

Today was Sol 11.

We undertook a morning EVA, the idea was the test the suits for the suitability whilst undertaking exploration of another planet. This was aimed at looking at oxygen in the blood after exercise, therefore the EVA was walking up to a ridge such as would be typical for explorations in the future.

Afterwards the crew members continued on with their individual’s projects. There was progress in processing the Eastern Veil Nebula photos, soy bacteria was grown in petri dishes this is for a logically first step identification method, in the radiation bacteria project 1 lettuce has sprouted and finally, the rover revival has continued where now 2 out of 4 wheels are functional, there is hope that it will make one EVA before the end of sim.

Due to poor weather most of the Crew stayed within the Hab today. Therefore, the group made cookies and prepped the pizza dough for Shannon’s arrival tomorrow.

Journalist Report – May 15th

Today was Sol 10.

This morning there was an EVA taken by two crew members who went to Marble ritual. The aim of the EVA was to continue on with drone testing. However, during the testing of the search and recover function with a dummy aeroplane, there was a take-off failure.

After the EVA, maintenance work was done on the Hab. The water pump was changed to prevent further leaking. After the installation the pressure increased enough to activate the hot water in the boiler. In addition the water filter was changed.

There was progress made on the processing of the all the photos of the pelican nebula and the robotic telescope finished the observations on the Western Veil Nebula. Poor weather prevented further observations in the solar telescopes. Work progresses on revising the rover, a final attempt will be made with a 2am conference tonight with the UK.

The crew continued on with their group problem solving game.

For outreach a space outfit was made for a Barbie to encourage women in STEM fields with a university in the UK.

Finally, the crew prepared homemade bread and pizza, harvesting tomatoes for both bruschetta and margarita pizza.

Journalist Report – May 14th

Good evening,

Today was Sol 9.

Today started with an engineering EVA to test the performance of the rovers; there have been issues with Spirits battery and Opportunities overheating. This involved traveling a total of 4 miles; the EVA was taken at 11am to try to access the rovers in the worst case scenario, midday heat. Curiosity performed fine, however Opportunity overheated after only 1.3miles and Spirits battery dramatically discharged.

In addition crew members cleaned and assessed the damage of the X-5 rover, the data from the black box was assessed to help with the next iteration of drone.

Another EVA was taken to establish contaminants in the sand, his was undertaken by Zubin’s head. The EVA whilst successful in terms of experimental point of view, there was difficulties with communicating with the Hab and due to a road being blocked the EVA was taken by walking, this involved an unfortunate oversight in timing and combined with the communication issues caused a panic in the Hab. This situation is detailed in the EVA report.

Kind Regards,

Zoe Townsend

journalist Report- May 02nd

Today was our eleventh simulation day. The commander and vice-commander began the day with a nice morning EVA along the Hab Ridge Road. The rest of the Crew woke up at 7:30 am and we had breakfast at 8:00.

After the breakfast, the commander and vice-commander deployed their respective experiment during an engineering EVA while the rest of the team finalized their own experiments.

The deployment was postponed to this morning as different minor bugs prevented us to do it yesterday evening.The GreenHab officer harvested the Spirulina that was growing since more than a week and started to dry it.

Before lunch, the crew worked on their experiments.

At 12 am, Maxime prepared the lunch. After meal, we wanted to discuss about Nathan’s experience and his martian constitution but Chloé had to work on her experiment.

At 5pm, Nathan, Maxime, Chloé and Julien went to EVA to Grey Moon. They came back to the station at 6:30pm.

After that, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. We will eat rice with tomatoes.

Like usual, we will do a crew meeting to speak about the day.

Tonight we will all observe the stars, enjoying the clear sky that the Utah offers.

Regards,

Journalist Report – May 13th

Today was Sol 8.

Today started with a relaxed morning, the crew members worked on their own projects. There was the continuation of work on processing the photographs, more progress was made to recover the software of the rover with another 2am conference with the UK, and another batch of bacteria was mixed with soil to re-inoculate the seeds, the previous equipment has been sterilised and the bacteria has been killed.

An EVA was taken today which was 2 ½ hours long, it was an eventful EVA that was mixed with an engineering EVA. Firstly, we delivered another Mars mouse to the release point, after we went to the experiment site and started the two experiments. The first was to test the oxygen levels in the suits; a circuit was completed to show the differences in the oxygen levels. However, due to the air holes and lack of sealing a new experiment must be proposed. The drone experiment continued, with the testing the X-5 and the Palantir. A failure occurred in the Palantir during piloting, the damage seems unrepairable. Unfortunately, the EVA finished with a failure of Opportunity. Opportunity had to be towed back to the Hab during and engineering EVA.